# Resilience for Regular Path Queries: Towards a Complexity Classification

### v2

@@ 7. Additional Complexity Results
Thus, this construction maintains the one-to-one correspondence between the contingency sets of $Q_{L}$ for $D$ and the cuts of the network that have finite cost: selecting facts to remove in the contingency set corresponds to cutting edges in the network, and ensuring that the contingency set does not satisfy the query corresponds to ensuring that no source to target paths remain. Furthermore, this the sum of multiplicities of the facts in a contingency set is equal to the cost of the corresponding cut, so that minimum cuts correspond to minimum contingency sets. We can then, as in the case of local languages, simply solve the MinCut problem for this network in $\tilde{O}(|A|\times|D|^{2}\times|\Sigma|^{2})$, concluding the proof. ∎
As for chain languages $L$ that are not bipartite, we conjecture that $\mathtt{RES}_{\mathrm{set}}(L)$ is always NP-hard. We leave this open, but show hardness for the specific non-bipartite chain language of Example 7.3:
{toappendix}
7.3. Proof of Section 7.3
{propositionrep}
The resilience problem for the language $ab|bc|ca$ is NP-hard.
Proof.
Figure 15. Gadget used in the proof of Section 7.3
We claim that the gadget shown in Figure 15 is valid for $ab|bc|ca$. Indeed, the gadget has the following properties: We can see that it meets all the criteria of being a pre-hardness gadget. We can also generate a condensed hypergraph of matches that forms an odd chain, and hence this gadget is a valid hardness gadget for $ab|bc|ca$. ∎
{toappendix}
7.4. Proof of Section 7.4: Tractability via Submodularity
Tractability due to submodularity.
We now present our second case of tractable languages. All PTIME cases discussed in this paper, as well as those discussed in previous work on resilience for conjunctive queries (DBLP:journals/pvldb/FreireGIM15; DBLP:conf/pods/FreireGIM20; makhija2024unified), are established by reducing to MinCut problems in a flow graph constructed in linear-time data complexity. But we now show that, in our setting, some languages are tractable for a different reason, by reducing to submodular function minimization. This implies a PTIME algorithm (MCCORMICK2005321), albeit with a worse polynomial degree than when reducing to MinCut.
{propositionrep}
[Tractability via submodularity] Let $\alpha=a_{1}\cdots a_{n}$ be a word where all letters are distinct, and let $a_{n+1}$ be a letter not occurring in $\alpha$. The problem $\mathtt{RES}_{\mathrm{bag}}(\alpha|a_{n-1}a_{n+1})$ is in PTIME.
Hence, the query $abc|be$, or the query $abcd|ce$ from Figure 1, enjoy tractable resilience.
{proofsketch}
We define a function on sets $Y$ of domain elements whose value is the resilience when we commit to removing precisely the $a_{n+1}$-facts whose tail is in $Y$. We show that this function is submodular, relying on a known result on the modularity of maximal flows (megiddo1974optimal, Lemma 3.2(ii)), which implies that we can find the set $Y$ minimizing the function and compute the resilience.
{toappendix}
We first recall the standard definition of submodular set functions:
Definition 7.6 (Submodular set function).
A submodular set function is a function of the form $f\colon 2^{V}\rightarrow\mathbb{R}$ that satisfies the following property: for every $A\subseteq S\subseteq V$ and $x\in S\setminus B$, we have that $f(A\cup\{x\})-f(A)\geq f(B\cup\{x\})-f(B)$.
Our result then hinges on a lemma about the resilience of singleton languages:
{lemmarep}
Let $\alpha=a_{1}\cdots a_{m}$ be a non-empty word without repeated letters, let $L=\{\alpha\}$ be the corresponding singleton language, and consider the query $L$. Let $D$ be a bag database (seen as a multiset). Then the function $f\colon 2^{\mathrm{Adom}(D)}\rightarrow\mathbb{N}$ which maps $Y\subseteq\mathrm{Adom}(D)$ to $\mathtt{RES}_{\mathrm{bag}}(L,D\setminus\{a_{m}(u,v)\in D\mid u\in Y\})$ is a submodular set function.
In other words, if $\alpha$ has no repeated letters, then $\mathtt{RES}_{\mathrm{bag}}(\{\alpha\})$ is a submodular function of the set $Y$ of domain elements on which we remove all outgoing $a_{m}$-edges. Let us prove the lemma, using a known result from the study of multi-target flow problems (megiddo1974optimal, Lemma 3.2(ii)).
Proof of Definition 7.6.
To prove this, we formulate $\mathtt{RES}_{\mathrm{bag}}(L)$ as a single-source multi-target flow problem and use the existing submodularity result of (megiddo1974optimal).
Since $L$ is a singleton language without repeated letters, we know through the proof of Theorem 3.3 that it can be encoded as a MinCut problem on a product network construction. We can modify the product network construction to have multiple target nodes, with one target for each element of $\mathrm{Adom}(D)$. More precisely, let $q_{0},\ldots,q_{m}$ be the successive states of the obvious DFA for the language $\{\alpha\}$. The network as previously constructed consists of one target node $t_{\text{target}}$, and edges with capacity $\infty$ from nodes of the form $(v,q_{m})$ to $t_{\text{target}}$, then for each fact $f=a_{m}(u,v)$ there is an edge of weight $\mathtt{mult}(D,f)$ from $(u,q_{m-1})$ to $(v,q_{m})$. The nodes of the form $(v,q_{m})$ have no other incoming or outgoing edges, and the nodes of the form $(u,q_{m-1})$ have no other outgoing edges.
It is then clear that we can modify the product network to be a multi-target network and to have targets corresponding to each element of $\mathrm{Adom}(D)$: for each $u\in\mathrm{Adom}(D)$, letting $\kappa$ be the total number of facts of the form $a_{m}(u,v)$ in the (bag) database $D$, replace the outgoing edges from $(u,q_{m-1})$ with one outgoing edge with capacity $\kappa$ going to a fresh target node $t_{u}$. (In particular, when there are no facts of the form $a_{m}(u,v)$, then the node $(u,q_{m-1})$ had no outgoing edges and we add a useless outgoing edge of capacity zero to a fresh target node $t_{u}$. It is clear that this does not change the value of the flow.
The construction ensures that there is a bijection between the set of targets $t_{u}$ of the modified flow graph and $\mathrm{Adom}(D)$. It is also clear that, for any $Y\subseteq\mathrm{Adom}(D)$, the resilience $\mathtt{RES}_{\mathrm{bag}}(L,D\setminus\{a_{m}(u,v)\in D\mid u\in Y\})$ is the value of the flow in the network where we remove the targets $\{t_{a}\mid a\in Y\}$. Now, for every subset $T$ of the targets $\{t_{a}\mid a\in\mathrm{Adom}(D)\}$, let $mc(T)$ be the cost of a minimum cut that separates the source from all targets of $T$. (In particular, the cost of a minimum cut of the original product network is $mc(C)$.) It is then known (megiddo1974optimal, Lemma 3.2(ii)), that for a flow graph with a single source and multiple targets, the function $mc(T)$ is a submodular set function. It is also known that if $f\colon 2^{V}\to\mathbb{N}$ is a submodular function then so is $g\colon 2^{V}\to\mathbb{N}$ defined by $g(A)\coloneq f(V\setminus A)$ (Clark2016, Lemma 1.5). Thus, the function that maps $Y\subseteq\mathrm{Adom}(D)$ to $mc(\mathrm{Adom}(D)\setminus Y)$ is also submodular. As we explained, $mc(\mathrm{Adom}(D)\setminus Y)$ is precisely the resilience of $L$ on $D$ with the facts of the form $\bigcup_{u\in Y}a_{m}(u,\_)$ removed. Thus, through the connection between the value of the flow network and the output to the resilience problem $\mathtt{RES}_{\mathrm{bag}}(L,D\setminus X)$, we can conclude that the desired function is submodular, as we claimed. ∎
We then use this lemma to show Section 7.4
Proof of Section 7.4.
Let $\alpha=a_{1}\cdots a_{n}$ and $\alpha^{\prime}=a_{n-1}a_{n+1}$ where all letters $a_{1},\ldots,a_{n+1}$ are distinct, and consider the finite reduced language $L=\alpha|\alpha^{\prime}$. The query $Q_{L}$ is equivalent to the following UCQ:
$\exists x_{1}\cdots x_{n+5}[\big{(}A_{1}(x_{1},x_{2})\wedge\cdots\wedge A_{n}(x_{n},x_{n+1})\big{)}\vee\big{(}A_{n-1}(x_{n+2},x_{n+3})\wedge A_{n+1}(x_{n+4},x_{n+5})\big{)}].$
For each value $v$ of the domain of the input database $D$, let $a_{i}(v,\_)$ be the multiset of edges of the form $\bullet\xrightarrow{a_{i}}v$ having $v$ as head, and similarly define $a_{i}(\_,v)$ to be the multiset of edges labeled $a_{i}$ with $v$ as source. We would like to express the function $\mathtt{RES}_{\mathrm{bag}}(\alpha|\alpha^{\prime},D)$ in terms of the function $\mathtt{RES}_{\mathrm{bag}}(\alpha)$ and apply Definition 7.6.
To do this, we first observe that in any contingency set $S$, for any value $v$ in the domain, we must have either all edges of the form $a_{n-1}(\_,v)$ in $S$ or all edges of the form $a_{n+1}(v,\_)$ in $S$. (In particular, if one of these two sets is empty then we can vacuously take in $S$ the empty set of its edges.) Indeed, we can see that all edges of the form $a_{n-1}(\_,v)$ and $a_{n+1}(v,\_)$ create an $a_{n-1}a_{n+1}$-walk. Thus, if both sets are not fully included, then the facts of the nonempty set $a_{n-1}(\_,v)\setminus S$ together with those of the nonempty set $a_{n+1}(v,\_)\setminus S$ create an $a_{n-1}a_{n+1}$-walk in $D\setminus S$, contradicting the fact that $S$ is a contingency set. So indeed, for each value $v\in\mathrm{Adom}(D)$, any contingency set must contain all edges of at least one of the two sets $a_{n-1}(\_,v)$ or $a_{n+1}(v,\_)$.
Now we can express the resilience of the language $\alpha|a_{n-1}a_{n+1}$ in terms of the resilience of the language $\alpha$. For each value $v$ which is the middle vertex of an $a_{n-1}(\_,v)a_{n+1}(v,\_)$-walk, we can decide whether to include all edges from $a_{n-1}(\_,v)$ or all edges from $a_{n+1}(v,\_)$ in the contingency set and compute the resilience of the query $\alpha$ over the database with these edges removed. This ensures that there can be no matches of the query disjunct $a_{n-1}a_{n+1}$. Thus, the resilience of the query $\alpha|a_{n-1}a_{n+1}$ can be computed as:
(1) $\mathtt{RES}_{\mathrm{bag}}(\alpha|\alpha^{\prime},D)=\min_{Z\subseteq\mathrm{Adom}(D)}\Big{\{}\sum_{v\in Z}|a_{n-1}(\_,v)|+\sum_{v\not\in Z}|a_{n+1}(v,\_)|+\mathtt{RES}_{\mathrm{bag}}(\alpha,D\setminus\bigcup_{v\in Z}a_{n-1}(\_,v))\Big{\}}$
where cardinalities denote multiset cardinalities. In other words, the resilience is computed by minimizing over subsets $Z$ of the domain elements: for $v\in Z$ we remove all incoming $a_{n-1}$-edges, for $v\notin Z$ we remove all outgoing $a_{n+1}$-edges, and then we find the resilience problem for the query $\alpha$ on the database where the $a_{n-1}$-edges leading into vertices of $Z$ have been removed.
Let $D_{n-1,Z}\coloneq D\setminus\bigcup_{v\in Z}a_{n-1}(\_,v))$, which occurs is the expression above, and let $D_{n,Z}\coloneq D\setminus\bigcup_{v\in Z}a_{n}(v,\_))$. We make the following claim, dubbed (*): $\mathtt{RES}_{\mathrm{bag}}(\alpha,D_{n-1,Z})=\mathtt{RES}_{\mathrm{bag}}(\alpha,D_{n,Z})$, i.e., the resilience of $D$ with all edges of form $a_{n-1}(\_,v)$ removed for all $v\in Z$ is the same as the resilience of $D$ with all edges of form $a_{n}(v,\_)$ removed for all $v\in Z$. Indeed, to show (*), first observe that minimum contingency sets for $\alpha$ on $D_{n-1,Z}$ never include edges of $a_{n}(v,\_)$ for $v\in Z$, because such edges never occur in $\alpha$-walks given that $v$ has no incoming $a_{n-1}$-edge; symmetrically, minimum contingency sets for $\alpha$ on $D_{n,Z}$ never include edges of $a_{n-1}(\_,v)$ for $v\in Z$, because such edges never occur in $\alpha$-walks given that $v$ has no outgoing $a_{n}$-edge. Now, to finish the proof of (*), letting $D^{\prime}\coloneq D\setminus\left(\bigcup_{v\in Z}a_{n-1}(\_,v)\cup\bigcup_{v\in Z}a_{n}(v,\_)\right)$, it suffices to show that any subset $S\subseteq D^{\prime}$ is a contigency set for $D_{n-1,Z}$ iff it is one for $D_{n,Z}$. Indeed, for one direction, taking such an $S$ such that $D_{n-1,Z}\setminus S$ does not satisfy $Q_{\alpha}$, it suffices to show that adding back the facts $\bigcup_{v\in Z}$ and removing those of $D\setminus\bigcup_{v\in Z}a_{n}(v,\_)$ cannot create a query match. But the added facts cannot be part of an $\alpha$-walk, because this walk would necessarily have $v\in Z$ as its last vertex, but there are no $a_{n}$-edges leaving out of $v$ in $D_{n,Z}$. The other direction is symmetrical. Hence, we have established (*).
From (*), we can thus restate Equation 1 as:
(2) $\mathtt{RES}_{\mathrm{bag}}(\alpha|\alpha^{\prime},D)=\min_{Z\subseteq\mathrm{Adom}(D)}\Big{\{}\sum_{v\in Z}|a_{n-1}(\_,v)|+\sum_{v\not\in Z}|a_{n+1}(v,\_)|+\mathtt{RES}_{\mathrm{bag}}(\alpha,D\setminus\bigcup_{v\in Z}a_{n}(\_,v))\Big{\}}$
We next show that Equation 2 is a submodular minimization problem, and can be solved in PTIME. In other words, having fixed the input (multiset) database $D$, for $\mathtt{RES}_{\mathrm{bag}}(\alpha|\alpha^{\prime},D)=\min_{Z\subseteq\mathrm{Adom}(D)}f(D,Z)$ with $f$ defined as the expression inside the minimization in Equation 2 above, we would like to show that $f(D,Z)$ is submodular. The sum of submodular functions is a submodular function, so it suffices to show that each term of the sum is a submodular function in $Z$.
We can show that the function $f_{1}(Z)=\sum_{v\in Z}|a_{n-1}(\_,v)|$ is submodular because for each $Z\subseteq\mathrm{dom}(D)$ and $v\in\mathrm{dom}(D)\setminus Z$, we have $f_{1}(Z\cup\{v\})-f_{1}(Z)=|a_{n-1}(\_,v)|$. Further, $f_{2}(Z)=\sum_{v\not\in Z}|a_{n+1}(v,\_)|$ is submodular for the same reason: for each $Z\subseteq\mathrm{dom}(D)$ and $v\in\mathrm{dom}(D)\setminus Z$, we have $f_{2}(Z\cup\{v\})-f_{2}(Z)=-|a_{n+1}(v,\_)|$. Finally, since $\alpha$ is a non-empty word without repeated letters, it follows from Definition 7.6 that $f_{3}(Z)=\mathtt{RES}_{\mathrm{bag}}(\alpha,D\setminus\bigcup_{v\in Z}a_{n}(v,\_))$ is submodular over $\mathrm{Adom}(D)$, since $\alpha$ is a chain query.
In this way, we have shown that the resilience of the query $\alpha|\alpha^{\prime}$ can be phrased as a submodular minimization problem, and hence can be computed in PTIME. ∎
Interestingly, this proof technique does not apply even if we add a single letter to right of the first word - in particular, as we pointed out in the main text, we do not know if the resilience of the query $abce|be$ can be computed in PTIME. We also do not know what happens when the second word is longer, i.e., we do not know if the resilience of the query $abc|bef$ can be computed in PTIME. Note that some variants of the query are four-legged (e.g., $abc|ebf$, which up to renaming is $axb|cxd$, for which resilience is NP-hard by Proposition 4.8), and others are not four-legged but covered by Section 7.5.
Additional NP-hard cases and open problems.
@@
Already in the case of finite reduced languages, there are gaps remaining between the tractable cases (local languages, and languages covered by Section 7.2 or Section 7.4) and the hard cases (languages containing a word with a repeated letter, and four-legged languages) that we have identified so far. We show below that some of these unclassified languages are NP-hard by building specific gadgets:
{toappendix}
7.5. Proof of Section 7.5
{propositionrep}
@@
Figure 16. Gadget used in the proof of Section 7.5 regarding $\mathtt{RES}_{\mathrm{set}}(abcd|be|ef)$
Figure 17. Gadget used in the proof of Section 7.5 regarding $\mathtt{RES}_{\mathrm{set}}(abcd|bef)$
We claim that the gadget shown in Figure 16 is valid for $abcd|be|ef$ and the gadget shown in Figure 17 is valid for $abcd|bef$. We can see that the gadgets meet all the criteria of a pre-hardness gadget, and show that there exists a condensed hypergraph of matches of the pre-gadgets for these queries that is an odd chain. These properties are checked programmatically, and we conclude that these gadgets are valid hardness gadgets for $abcd|be|ef$ and $abcd|bef$. ∎
We leave the question open for other languages, including $abcd|be$, $abc|bcd$, or $abc|bef$. Many infinite languages also remain unclassified, such as $ax^{*}b|xd$ for instance.

### v3

@@ 7.1. Tractability of Bipartite Chain Languages
Thus, this construction maintains the one-to-one correspondence between the contingency sets of $D$ for $Q_{L}$ and the cuts of the network that have finite cost: selecting facts to remove in the contingency set corresponds to cutting edges in the network, and ensuring that the contingency set does not satisfy the query corresponds to ensuring that no source to target paths remain. Furthermore, this is the sum of multiplicities of the facts in a contingency set is equal to the cost of the corresponding cut, so that minimum cuts correspond to minimum contingency sets. We can then, as in the case of local languages, simply solve the MinCut problem for this network in $\tilde{O}(|A|\times|D|^{2}\times|\Sigma|^{2})$, concluding the proof. ∎
7.2. One-Dangling Languages
We now present another example of tractable languages, dubbed one-dangling languages, for which tractability is shown by a more involved reduction to a flow problem.
Definition 7.8 (One-dangling language).
A one-dangling language is a language that can be written as $L\cup\{xy\}$ where $L$ is a local language over some alphabet $\Sigma$ and where $x$ and $y$ are distinct letters such that at least one of them is not in $\Sigma$.
Note that one-dangling languages cover in particular the more restricted class of languages for which tractability had been shown in the conference version of this paper (amarilli2025resilience, , Proposition 7.7). They also cover the two languages $abcd|be$ and $ax^{*}b|xd$ for which the complexity of resilience had been left open in (amarilli2025resilience, ). So we can subsume that result by showing the following:
Proposition 7.9.
[…]
In addition to covering more languages, this result is a combined tractability result instead of a data complexity result, and it gives a better polynomial exponent than (amarilli2025resilience, , Proposition 7.7). This is because our proof will proceed via a reduction to a flow problem, instead of leveraging the more complicated technique of submodular function minimization used in (MCCORMICK2005321, ).
We prove Proposition 7.9 in the rest of this subsection, before finally turning to other NP-hard cases in the rest of the section. We note that an argument used the proof was first posted online (cstheory_martin, ). We do a slight abuse of notation throughout the proof for convenience: for $D^{\prime\prime}$ a bag database and $L^{\prime\prime}$ a language, we will talk of a contingency set of $D^{\prime\prime}$ for $L^{\prime\prime}$ simply to mean a contingency set of $D^{\prime\prime}$ for $Q_{L^{\prime\prime}}$.
Due to Proposition 6.3, if there is a PTIME solution for $\mathtt{RES}_{\mathrm{bag}}(L\cup\{xy\})$, then there is a PTIME solution for $\mathtt{RES}_{\mathrm{bag}}(L^{R}\cup\{yx\})$, and the mirror of a local language is local (as can easily be seen from Proposition 3.5). Hence, we can assume without loss of generality that $y\not\in\Sigma$.
[…]
Let us now define how the input bag database $D$ over $\Sigma\cup\{y\}$ is rewritten to a bag database $D^{\prime}$ over $\Sigma\cup\{z\}$. To do this, we will extend the semantics of bag graph databases to what we call the extended bag semantics, where we allow the function $\mathtt{mult}$ to give negative multiplicities, or multiplicity zero, to edges of the database. The rest of the semantics remains unchanged, and the definition of resilience is unchanged as well – noting that, when computing resilience over database in extended bag semantics, one can always assume that facts with a non-positive multiplicity are removed. For this reason, it is easy to see that $\mathtt{RES}_{\mathrm{bag}}(L)$ under the extended semantics reduces in linear time to $\mathtt{RES}_{\mathrm{bag}}(L)$ under the normal bag semantics. Let us use the notation $\mathtt{RES}_{\mathrm{bag}}^{\sf ex}(L)$ for a fixed language $L$ to denote the resilience problem when posed over databases in extended bag semantics.
We can now describe the transformation. Given a bag database $D$, let $V$ be the domain of $D$. For every node $v\in V$, let $D^{-}_{x}(v)$ be the set of edges of the form $u\xrightarrow{x}v$ for some $u\in V$, and let $D^{+}_{y}(v)$ be the set of edges of the form $v\xrightarrow{y}u$ for some $u\in V$. We rewrite $D$ by doing the following for each $v\in V$:
  1. (1)
    add a fresh node $(v,{\sf in})$ into $V$;
  2. (2)
    replace all edges of the form $u\xrightarrow{x}v\in D^{-}_{x}(v)$ by the edge $u\xrightarrow{x}(v,{\sf in})$, keeping the same multiplicity;
  3. (3)
    add an edge $(v,{\sf in})\xrightarrow{z}v$ into $D$ and define $\mathtt{mult}((v,{\sf in})\xrightarrow{z}v)\coloneq\sum_{e\in D^{-}_{x}(v)}\mathtt{mult}(e)-\sum_{e\in D^{+}_{y}(v)}\mathtt{mult}(e)$ (which may be negative); and
  4. (4)
    erase any edges in $D^{+}_{y}(v)$ from $D$.
Let us call $D^{\prime}$ the resulting bag database: it is a bag database in extended semantics, and it is over the alphabet $\Sigma\cup\{z\}$ because applying the last point of the list above over all $v\in V$ will remove all $y$-facts. Let us define the value $\kappa\coloneq\sum_{v\in V}\sum_{e\in D^{+}_{y}(v)}\mathtt{mult}(e)$ (evaluating this expression in $D$). Note that $\kappa$ can be computed in linear time from $D$ as it is simply the sum of multiplicities of all edges labeled $y$ in $D$.
We now claim the following:
[…]
which suffices to conclude, because we have computed $\kappa$ and we can compute the resilience expression in the right-hand side using Theorem 3.13: the language $L^{\prime}$ is local, and as we explained the complexity of computing resilience is unchanged when working over extended bag semantics databases. To show Equation ($\dagger$7.2), we will show the following claim, and explain after its proof how the lemma implies Equation ($\dagger$7.2):
Claim 7.10.
(i) For any minimal contingency set $C$ of $D$ for $L\cup\{xy\}$, there exists a contingency set $C^{\prime}$ of $D^{\prime}$ for $L^{\prime}$ such that $\sum_{e\in C}\mathtt{mult}(e)=\kappa+\sum_{e\in C^{\prime}}\mathtt{mult}(e)$; and (ii) for any minimal contingency set $C^{\prime}$ of $D^{\prime}$ for $L^{\prime}$, there exists a contingency set $C$ of $D$ for $L\cup\{xy\}$ such that $\sum_{e\in C}\mathtt{mult}(e)=\kappa+\sum_{e\in C^{\prime}}\mathtt{mult}(e)$.
We end the proof of Proposition 7.9 by explaining why Equation ($\dagger$7.2) follows from 7.10. First, suppose that $\mathtt{RES}_{\mathrm{bag}}(L\cup\{xy\},D)<\mathtt{RES}_{\mathrm{bag}}^{\sf ex}(L^{\prime},D^{\prime})+\kappa$. Let $C^{*}$ be a minimal contingency set of $D$ for $L\cup\{xy\}$, and let $C^{\prime*}$ be a minimal contingency set of $D^{\prime}$ for $L^{\prime}$. We have that $\sum_{e\in C^{*}}\mathtt{mult}(e)<\kappa+\sum_{e\in C^{\prime*}}\mathtt{mult}(e)$. However, due the to point (i) in 7.10, we know that there exists a contingency set $C^{\prime}$ of $D$ for $L^{\prime}$ such that $\sum_{e\in C^{*}}\mathtt{mult}(e)=\kappa+\sum_{e\in C^{\prime}}\mathtt{mult}(e)$, and so $\sum_{e\in C^{\prime}}\mathtt{mult}(e)<\sum_{e\in C^{\prime*}}\mathtt{mult}(e)$, which contradicts the fact that $C^{\prime*}$ is minimal. Now suppose that $\mathtt{RES}_{\mathrm{bag}}(L\cup\{xy\},D)>\mathtt{RES}_{\mathrm{bag}}^{\sf ex}(L^{\prime},D^{\prime})+\kappa$. We reach a contradiction using an analogous argument and using point (ii) in 7.10. Thus, we have concluded the proof of Proposition 7.9.
7.3. Additional NP-Hard Cases
@@
Already in the case of finite infix-free languages, there are gaps remaining between the tractable cases (local languages, and languages covered by Proposition 7.6) and the hard cases (languages containing a word with a repeated letter, and four-legged languages) that we have identified so far. We show below that an unclassified languages is NP-hard by building a specific gadget:
Proposition 7.11.
@@
Figure 15. Completed gadget used in the proof of Proposition 7.11 regarding $\mathtt{RES}_{\mathrm{set}}(abcd|be|ef)$
Figure 16. Completed gadget used in the proof of Proposition 7.11 regarding $\mathtt{RES}_{\mathrm{set}}(abcd|bef)$
We claim that the gadget shown in Figure 15 is valid for $abcd|be|ef$ and the gadget shown in Figure 16 is valid for $abcd|bef$ (also, they happen to be identical). We can see that the gadgets meet all the criteria of a pre-gadget, and show that there exists a condensed hypergraph of matches of the pre-gadgets for these queries that is an odd path. These properties can be checked by exhaustively considering all walks in the gadget, and we conclude that these gadgets are valid hardness gadgets for $abcd|be|ef$ and $abcd|bef$. ∎